In an experiment designed to measure the speed of light, a laser is aimed at a mirror that is 50 km due north. A detector is placed 100 m due east of the laser. The mirror is to be aligned so that light from the laser reflects into the detector. When properly aligned, what angle (in degree) should the normal to the surface of the mirror make with due south?



1
Expert's answer
2021-11-04T16:20:01-0400

The arrangement forms a right angled triangle.

Now angle of incidence = angle of reflection( law of reflection).

"tan\theta=\frac{100}{50\times 1000}=100/50000=0.002"

"tan^{-1} 0.002=0.1146^0"

Required angle "=00.1146\/2=0.0573^0"